Im wondering what the answer for x+3 over 8 = x over 2 is and if its open, false, idenity, or no solution. I tried multiplying the 8 on both sides first but then i got stuck on how to do the x+3 and the rest. If you have any help that would be great, thank you(:
Answer by rfer(16322)

(x+3)/8=x/2
8x=2(x+3)
8x=2x+3
6x=3
x=3/6
x=1/2